# Topological phase transition through tunable nearest-neighbor interactions in a one-dimensional lattice

### v1

@@ III.1 Phase diagram at half-filling
$S_{\text{BO}}(k)=\frac{1}{L^{2}}\sum_{i,j}e^{ik|i-j|}\langle B_{i}B_{j}\rangle,$ (11)
and
$S(k)=\frac{1}{L^{2}}\sum_{i,j}e^{ik|i-j|}\langle n_{i}n_{j}\rangle,$ (12)
where, $S_{\text{BO}}(k)$ is the bond order structure factor and $S(k)$ is the charge-density wave structure factor. To minimize the boundary effects, we consider bond and density correlation functions in the range from $L/4$ to $43L/$ 4 throughout our calculations. The extrapolated values of $S_{\text{BO}}(k=\pi)$ and $S(k=\pi/2)$ are shown in Fig. 6 as black diamonds and green stars respectively as a function of $V$ (along the dashed line in Fig. 2). The finite values of $S(k=\pi/2)$ for $V\lesssim-4.3$ and $V\gtrsim 4.3$ clearly indicates the CDW-II regions. At the same time, we find that the $S_{\text{BO}}(k=\pi)$ is finite in the region $-4.3<V<4.3$ except at $V=0$ where the system is in the SF phase. However, beyond $V\sim\pm 4.3$, the decrease in $S_{\text{BO}}(k=\pi)$ to smaller values and a sharp increase in $S(k=\pi/2)$ clearly indicates a transition to the CDW-II phase.
Figure 6: Extrapolated values of $S_{\text{BO}}(\pi)$ (black diamonds) and $S(\pi/2)$ (green stars) as a function of $V$ computed by taking a maximum system size of $L=200$.
To accurately determine the phase transition boundary between the BO and the CDW-II phases, we perform a finite-size scaling of the CDW-II structure factor $S(k=\pi/2)$. Our results show that the BO to CDW-II phase transition belongs to the Ising universality class. The critical point is identified through the appropriate scaling behavior of the CDW-II structure factor
$S(\pi/2)L^{2\beta/\nu}=F[(V_{1}-V_{1}^{c})L^{1/\nu}],$ (13)
where $\beta$ and $\nu$ are the magnetization and correlation length critical exponents Mondal et al. (2022a); Mishra et al. (2011). We plot $S(\pi/2)L^{2\beta/\nu}$ as a function of $V_{1}$ for $V_{2}=0$ in Fig. 7 (a) with $\beta=1/8$ and $\nu=1$. The curves for different system sizes are found to cross at the critical point $V_{1}^{c}\sim-4.786$. Moreover, a perfect data collapse for various system sizes is achieved using the scaling function in Eq. (13) at $V_{1}^{c}=-4.786$ as shown in Fig. 7 (b) which confirms the critical point. This approach allows us to map out the entire phase boundary as presented in the upper left part of Fig. 2 (a) (shown as black diamonds). A similar situation is found upon varying $V_{2}$ for each values of $V_{1}$ resulting in the phase boundary shown in the lower right part of Fig. 2 (a).
The upper left (lower right) part of the phase diagram in Fig. 2 (a) shows that as the value of $V_{2}$ ($V_{1})$ is increased, the critical point of the BO to CDW-II transition slowly shifts towards smaller values of attractive $V_{1}$ ($V_{2}$). Moreover, the transition boundary seems to saturate at a particular value of $V_{1}$ ($V_{2}$) in the limit of large $V_{2}$ ($V_{1}$).
Figure 7: (a) The figure illustrates the finite-size scaling of $S(\pi/2)$ according to Eq. 13 for various values of $V_{1}$ at fixed $V_{2}=0$. The intersection of the curves marks the critical point for the BO to CDW-II phase transition. (b) The scaled $S(\pi/2)$ is plotted against the scaled $V_{1}$, showing a data collapse onto a single curve, confirming $V_{1}^{c}=-4.786$ as the transition point.
The boundary in the limit of large $V_{2}$ ($V_{1}$) can be found using second-order perturbation theory, as shown in Appendix A. For example we find that for $V_{2}\gg t$ and $|V_{1}|$, the location of the boundary between the BO and the CDW-II phases in upper left part of Fig. 2 is given as
@@ III.2 Topological Properties
Figure 8: On-site particle density $\langle n_{i}\rangle$ plotted against the site index $i$ in the BOπ and BO0 phases at $V=-3.0$ (red circles) and $V=3.0$ (blue diamonds), respectively for a system size of $L=200$ sites.
One of the most significant consequences of bulk topology is the presence of degenerate edge modes in OBC. Therefore, as an initial step to characterize the topology of the system, we plot the on-site particle density $\langle n_{i}\rangle$ as a function of the site index for two different BO regimes. Fig. 8 shows the plot of $\langle n_{i}\rangle$ as a function of $i$ for $V=-3.0$ (red circles) and $V=3.0$ (blue diamonds) for a system size $L=200$. The larger values of $\langle n_{i}\rangle$ near the edge sites for $V=-3.0$ compared to uniform density distribution for $V=3.0$ suggests a topological difference between the two types of BO phases.
To concretely establish the topological nature of the BO phase appearing in the regime $-4.3\lesssim V\lesssim 0$ and to characterize the topological phase transition between the two BO phases at $V=0$, we calculate the bulk topological invariant of the system in this regime at various values of $V$ for a system size of $L=16$ using ED. The topological invariant for the interacting system is calculated using twisted boundary conditions, where the hopping strength at the boundary is modified as $t\xrightarrow{}te^{i\theta}$, with $\theta$ being the twist angle Zak (1989); Resta and Sorella (1995). As $\theta$ is varied from 00 to $2\pi$, the ground state wave function $|\psi(\theta)\rangle$ accumulates a phase, known as the Berry phase ($\gamma$), which is determined using the formula
@@
The Berry phase is 00 if the bulk of the system is trivial whereas it is $\pi$ if the bulk is topological in nature. We compute $\gamma$ for different values of $V$ and plot $\gamma/\pi$ as a function of $V$ in Fig. 9(a) (blue diamonds). The plot shows that $\gamma/\pi$ changes from $1$ to 00 at $V=0$ signifying a phase transition from a topological to a trivial BO phase at $V=0$. Based on the distinct Berry phases of the two BO phases, we denote the topological and the trivial BO phases as BOπ and BO0 respectively (Fig. 2 (a)).
@@
where $r=|i-j|$ and $Z(i)=1-2a^{\dagger}_{i}a_{i}$. We consider $i=2$ and $k=L-1$ to avoid the edge sites. We calculate $O_{\text{str}}$ for a system with $L=200$ sites and plot in Fig. 9(a) denoted as red stars. It is clear that before $V=0$, the string order parameter is finite and vanishes after $V=0$ which agrees well with the topological invariant (blue diamonds).
Figure 9: (a) Plots of the winding number $\gamma/\pi$ (blue diamonds) and the string order parameter $O_{\text{str}}$ (red stars) as a function of $V$. The density-density correlation matrix ($\Gamma$) for $V=-1.0$ and $1.0$ shown in (b) and (c), respectively, are calculated using the ED method for a system size of $L=16$ sites.
@@
We plot $\Gamma_{i,j}$ for $V=-1.0$ and $V=1.0$ in Figs. 9 (b) and (c) respectively. The presence of finite correlation matrix elements at the two corners, i.e., at the two isolated blue regions at opposite ends of the correlation matrix shown in Fig. 9 (b) indicates the existence of edge states for $V=-1.0$. This clearly signifies the topological character of the BOπ phase. In contrast, Fig. 9 (c) shows no such edge states, indicating the trivial nature of the BO0 phase for $V=1.0$.
@@
where the $\Lambda_{i}$’s correspond to the singular values at the central bond of the system in the MPS wave function of the ground state. We plot the lowest four values of the entanglement spectrum $\xi$ as a function of $V$ in Fig. 10 for a system size of $L=200$. We find a two-fold degeneracy in the region $-4.3\lesssim V<0$ apart from which the degeneracy is lifted for all the values of $V$. This indicates the topological nature of the BO ground state in the region $-4.3\lesssim V<0$.
Figure 10: Plot showing the lowest four values of the entanglement spectrum ($\xi_{1},~{}\xi_{2},~{}\xi_{3},~{}\xi_{4}$) as a function of $V$, for $L=200$.
@@
Figure 11: Evolution of the polarization $P(\tau)$ as a function of $\tau$ for a system with $L=200$ sites, shown for three distinct pumping cycles (1, 2, 3) corresponding to different values of $V_{c}$ ($=-0.8,~{}0.0,~{}1.0$) as indicated in the inset. In all three cases, we set $A_{\delta}=0.6~{}\text{and}~{}A_{\Delta}=0.6$. Blue diamonds, green squares, and red circles represent the polarization changes for cycles 1, 2, and 3, respectively. Robust charge pumping occurs in cycle 2, as it encloses the topological phase transition point $V=0.0$. In contrast, pumping breaks down (since there is a discontinuity in the polarization at $\tau=0.5$) in cycle 1, whereas no charge is pumped during cycle 3.
@@
Here we propose a quantity that dynamically probes and characterizes the aforementioned topological phase called Thouless charge pumping or topological charge pumping (TCP) Thouless (1983). TCP reveals a profound connection between topology and quantum transport, describing the quantized transport of particles through a cyclic and periodic modulation of system parameters without any external bias. The system parameters include one that drives the topological phase transition and an additional inversion symmetry-breaking term, ensuring that the bulk gap remains open throughout the pumping cycle — a necessary condition for particle pumping. The pumping cycle involves the adiabatic evolution of the system from a topological phase to a trivial phase and subsequently back to the initial topological phase again, with the cyclic path encircling the topological phase transition point. Due to the topological nature of the pump, an integer number of particles are transported across the system, and the number of particles transported is directly linked to the topological invariant of the system. Though originally proposed in the context of non-interacting systems, TCP has now been studied in systems of interacting particles. Several theoretical and experimental studies have adopted TCP as one of the crucial tools for revealing the signature of topological phases in both non-interacting as well as interacting systems Lohse et al. (2016); Nakajima et al. (2016); Citro and Aidelsburger (2023); Hayward et al. (2018); Rice and Mele (1982); Citro and Aidelsburger (2023); Asbóth et al. (2016b); Lin et al. (2020); Ke et al. (2017); Argüello-Luengo et al. (2024b); Kuno et al. (2017); Bertok et al. (2022); Mondal et al. (2022b); Hayward et al. (2018); Schweizer et al. (2016); Kraus et al. (2012); Taddia et al. (2017); Wang et al. (2013); Walter et al. (2023); Ke et al. (2020); Padhan et al. (2024); Parida et al. (2024); Jürgensen et al. (2023, 2021). In one dimension, the pumping protocol is given by the celebrated Rice-Mele model. For our system, the corresponding Rice-Mele model is written by adding a symmetry-breaking term, $\Delta$, defining our pumping Hamiltonian, which can now be expressed as
[…]
$+~{}(V_{c}-\delta(\tau)\sum\limits_{i}(-1)^{i}(n_{i}-\frac{1}{2})(n_{i+1}-\frac{1}{2})$
$+~{}\Delta(\tau)\sum\limits_{i}(-1)^{i}n_{i},$
@@
where $\tau$ is the pumping parameter. The parameter $\Delta$ ensures that the bulk gap remains open throughout the pumping cycle. $V_{c}$ is the origin of the pumping cycle in the $\delta-\Delta$ plane. $\delta(\tau)=A_{\delta}\text{cos}(2\pi\tau)$ and $\Delta(\tau)=A_{\Delta}\text{sin}(2\pi\tau)$ periodically modulate the nearest-neighbor interaction and the staggered potential term respectively. We consider three pumping cycles corresponding to three different values of $V_{c}$ as shown in the inset of Fig. 11. The parameter sets ($V_{c},A_{\delta},A_{\Delta}$) for these three cycles are taken to be ($-0.8t,0.6t,0.6t$), ($0.0t,0.6t,0.6t$) and ($1.0t,0.6t,0.6t$) for cycle 1,2, and 3 respectively. One should expect that pumping can only happen for cycle $2$ (green continuous line) as it winds around the topological phase transition point $V=0$. To study the TCP, we calculate a quantity called polarization given by
@@
where $i_{0}=(L+1)/2$ for a system of $L=200$ sites for all the three cycles. Here $\psi(\tau)$ represents the ground state wave function of the Hamiltonian defined in Eq. 19. The amount of charge pumped ($Q$) in a cycle is given as
[…]
Fig. 11 clearly demonstrates a smooth change in the polarization from $0.5$ to $-0.5$, indicating a total charge transfer of $|Q|=1$ for cycle 2. This signifies robust and quantized pumping occurring in cycle 2 as it encloses the gap-closing topological phase transition point as mentioned earlier. In contrast, cycle 1 exhibits a breakdown of pumping (due to the appearance of a discontinuity in the polarization), while cycle 3 shows no pumping at all Hayward et al. (2018). This analysis complements the topological phase transition between the two BO phases shown in the phase diagram [Fig. 2(a)].

### v2

@@ III.1 Phase diagram at half-filling
[…]
$\text{S}(\pi/2)L^{2\beta/\nu}=F[(V_{1}-V_{1}^{c})L^{1/\nu}],$ (12)
[…]
The upper left (lower right) part of the phase diagram in Fig. 2 (a) shows that as the value of $V_{2}$ ($V_{1})$ is increased, the critical point of the BO to CDW-II transition slowly shifts towards smaller values of attractive $V_{1}$ ($V_{2}$). Moreover, the transition boundary seems to saturate at a particular value of $V_{1}$ ($V_{2}$) in the limit of large $V_{2}$ ($V_{1}$). The boundary in the limit of large $V_{2}$ (or large $V_{1}$) can be found using second-order perturbation theory, as shown in Appendix A. For example we find that for $V_{2}\gg t$ and $|V_{1}|$, the location of the boundary between the BO and the CDW-II phases in upper left part of Fig. 2 is given as
@@ III.2 Topological Properties
Figure 10: On-site particle density $\langle n_{i}\rangle$ is plotted against the site index $i$ in the BOπ and BO0 phases at V=3.0V=-3.0 (red circles) and $V=3.0$ (blue diamonds), respectively for a system size of $L=200$ sites.
One of the most significant consequences of the bulk topology is the presence of degenerate edge modes in OBC. Therefore, as an initial step to characterize the topology of the system, we plot the on-site particle density $\langle n_{i}\rangle$ as a function of the site index for two different BO regimes. Fig. 10 shows the plot of $\langle n_{i}\rangle$ as a function of $i$ for V=3.0V=-3.0 (red circles) and $V=3.0$ (blue diamonds) for a system size $L=200$. The larger values of $\langle n_{i}\rangle$ near the edge sites for V=3.0V=-3.0 compared to uniform density distribution for $V=3.0$ suggests a topological difference between the two types of BO phases.
@@
$\gamma=i\int_{0}^{2\pi}d\theta~\langle\psi(\theta)|\partial_{\theta}|\psi(\theta)\rangle.$ (14)
The Berry phase is 0 if the bulk of the system is trivial whereas it is $\pi$ if the bulk is topological in nature. We compute $\gamma$ for different values of $V$ and plot $\gamma/\pi$ as a function of $V$ in Fig. 11(a) (blue diamonds). The plot shows that $\gamma/\pi$ changes from $1$ to 0 at $V=0$ signifying a phase transition from a topological to a trivial BO phase at $V=0$. Based on the distinct Berry phases, we denote the topological and the trivial BO phases as BOπ and BO0 respectively (Fig. 2 (a)). Additionally, note that the value of $\gamma/\pi=1$ in the CDW-II phase appearing in the region of $V<0$, despite its non-topological nature. However, this phase lacks other topological signatures, such as the presence of degenerate zero-energy edge states as discussed in Fig. 3. We support this conclusion through additional diagnostics discussed below, which collectively confirm the non-topological nature of the CDW-II phase.
The apparent nontrivial value of $\gamma/\pi$ in the CDW-II phase arises due to the finite-size system under consideration. Note that the CDW-II phase emerges due to the spontaneous inversion symmetry breaking, where the excitation gap remains finite for finite systems and closes only in the thermodynamic limit. However, one can see that the value of $G$ remains finite at the BOπ to CDW-II phase transition point (see Fig. 3 ). It is important to emphasize that a quantum phase transition requires the excitation gap $G_{n}=E_{1}(N)-E_{0}(N)$ to vanish, where $E_{0}(N)$ and $E_{1}(N)$ represent the ground state and first excited state energy, respectively Sachdev (2011). However, in this scenario, $G_{n}$ vanishes throughout the entire range of $V$ because, the BOπ phase is two-fold degenerate under OBC due to the presence of the edge states and the CDW-II phase is also two-fold degenerate. In such a situation, the BOπ and the CDW-II phases can not be distinguished from the excitation gap. In contrast, by employing PBC removes these degeneracies and reveals the intrinsic bulk gap of the BOπ phase. We plot the extrapolated value of $G_{n}$ in Fig. 11(b) under PBC which shows that the excitation gap closes at $V\sim-4.3$, confirming a phase transition from the BOπ to the CDW-II phase. Beyond this point ($V\lesssim-4.3$), the gap vanishes in the entire CDW-II phase, consistent with the expected degeneracy of the two $Z_{2}$-symmetry-broken ground states in the thermodynamic limit. Since a change in the topological invariant requires gap closing, the Berry phase remains unchanged across the transition in finite-size systems. This robustness of the topological invariant against spontaneous symmetry-breaking has been reported in earlier works Raj et al. (2021), and similar behavior has also been observed in related models Melo et al. (2023).
Figure 11: (a) Plot of the winding number $\gamma/\pi$ (blue diamonds) and the string order parameter $O_{\text{str}}$ (red stars) as a function of $V$. (b) Plot of the extrapolated value of $G_{n}$ under periodic boundary conditions, taking a maximum system size of $L=120$ sites. The vertical dashed lines in (a) separate different phases. The density-density correlation matrix $\Gamma$ for V=1.0V=-1.0 and $1.0$ is shown in (c) and (d), respectively, calculated using ED for a system size of $L=16$ sites.
@@
where r=|ik|r=|i-k| and $Z(i)=1-2a^{\dagger}_{i}a_{i}$. We consider $i=2$ and k=L1k=L-1 to avoid the edge sites. We calculate $O_{\text{str}}$ for a system with $L=200$ sites and plot in Fig. 11(a) denoted as red stars. It is clear that before $V=0$, the string order parameter is finite and vanishes after $V=0$ which agrees well with the topological invariant (blue diamonds). However, the string order parameter remains finite in the CDW-II phases Dalla Torre et al. (2006); Rossini and Fazio (2012); Batrouni et al. (2013) with opposite signs in the two regimes of interaction $V$. The change of sign is a consequence of two differing spatial arrangements of particles in the CDW-II phases for $V>0$ (i.e., $01100110$) and $V<0$ (i.e., $11001100$).
@@
We plot $\Gamma_{i,j}$ for V=1.0V=-1.0 and $V=1.0$ in Figs. 11 (c) and (d) respectively. The presence of finite correlation matrix elements at the two corners, i.e., the two isolated blue regions at opposite ends of the correlation matrix shown in Fig. 11 (c) indicates the existence of edge states for V=1.0V=-1.0 and hence the topological nature of the BOπ phase. In contrast, Fig. 11 (d) shows no such edge states, indicating the trivial nature of the BO0 phase for $V=1.0$.
@@
where the $\Lambda_{i}$’s correspond to the singular values at the central bond of the system in the MPS wave function of the ground state. We plot the lowest four values of the entanglement spectrum $\xi$ as a function of $V$ in Fig. 12 for a system size of $L=200$. We find a two-fold degeneracy in the region $4.3\lesssim V<0$ apart from which the degeneracy is lifted for all the values of $V$. This indicates the topological nature of the BO ground state in the region $4.3\lesssim V<0$. This diagnostic also complements the non-topological nature of the CDW-II phase arising for the $V\lesssim-4.3$ regime, claimed earlier.
Figure 12: The lowest four values of the entanglement spectrum ($\xi_{1},~\xi_{2},~\xi_{3},~\xi_{4}$) as a function of $V$, for $L=200$.
@@
Figure 13: Evolution of the polarization $P(\tau)$ as a function of $\tau$ for a system with $L=200$ sites, shown for three distinct pumping cycles (1, 2, 3) corresponding to different values of $V_{c}$ (=0.8,0.0,1.0=-0.8,~0.0,~1.0) as indicated in the inset. In all three cases, we set $A_{\delta}=0.6~\text{and}~A_{\Delta}=0.6$. Blue diamonds, green squares, and red circles represent the polarization changes for cycles 1, 2, and 3, respectively. Robust charge pumping occurs in cycle 2, as it encloses the topological phase transition point $V=0.0$. In contrast, pumping breaks down (since there is a discontinuity in the polarization at $\tau=0.5$) in cycle 1, whereas no charge is pumped during cycle 3.
@@
Here we propose a quantity that dynamically probes and characterizes the aforementioned topological phase called Thouless charge pumping or topological charge pumping (TCP) Thouless (1983). TCP reveals a profound connection between topology and quantum transport, describing the quantized transport of particles through a cyclic and periodic modulation of system parameters without any external bias. The system parameters include one that drives the topological phase transition and an additional inversion symmetry-breaking term, ensuring that the bulk gap remains open throughout the pumping cycle — a necessary condition for particle pumping. The pumping cycle involves the adiabatic evolution of the system from a topological phase to a trivial phase and subsequently back to the initial topological phase again, with the cyclic path encircling the topological phase transition point. Due to the topological nature of the pump, an integer number of particles are transported across the system, and the number of particles transported is directly linked to the topological invariant of the system. Though originally proposed in the context of non-interacting systems, TCP has now been studied in systems of interacting particles. Several theoretical and experimental studies have adopted TCP as one of the crucial tools for revealing the signature of topological phases in both non-interacting as well as interacting systems Lohse et al. (2016); Nakajima et al. (2016); Citro and Aidelsburger (2023); Hayward et al. (2018); Rice and Mele (1982); Citro and Aidelsburger (2023); Asbóth et al. (2016b); Lin et al. (2020); Ke et al. (2017); Argüello-Luengo et al. (2024); Kuno et al. (2017); Bertok et al. (2022); Mondal et al. (2022b); Hayward et al. (2018); Schweizer et al. (2016); Kraus et al. (2012); Taddia et al. (2017); Wang et al. (2013); Walter et al. (2023); Ke et al. (2020); Padhan et al. (2024); Parida et al. (2024); Jürgensen et al. (2023, 2021); Viebahn et al. (2024); Zhu et al. (2024). In one dimension, the pumping protocol is given by the celebrated Rice-Mele model. For our system, the corresponding Rice-Mele model is written by adding a symmetry-breaking term, $\Delta$, defining our pumping Hamiltonian, which can now be expressed as
$H(\tau)$ $=$ $-~t\sum\limits_{i}({a}^{\dagger}_{i}{a}_{i+1}+\text{H.c.})$ (18)
$+~(V_{c}-\delta(\tau)\sum\limits_{i}(-1)^{i+1}(n_{i}-\frac{1}{2})(n_{i+1}-\frac{1}{2})$
$+~\Delta(\tau)\sum\limits_{i}(-1)^{i+1}n_{i},$
@@
where $\tau$ is the pumping parameter. The parameter $\Delta$ ensures that the bulk gap remains open throughout the pumping cycle. $V_{c}$ is the origin of the pumping cycle in the $\delta-\Delta$ plane. $\delta(\tau)=A_{\delta}\text{cos}(2\pi\tau)$ and $\Delta(\tau)=A_{\Delta}\text{sin}(2\pi\tau)$ periodically modulate the NN interaction and the staggered potential term respectively. We consider three pumping cycles corresponding to three different values of $V_{c}$ as shown in the inset of Fig. 13. The parameter sets ($V_{c},A_{\delta},A_{\Delta}$) for these three cycles are taken to be (0.8t,0.6t,0.6t-0.8t,0.6t,0.6t), ($0.0t,0.6t,0.6t$) and ($1.0t,0.6t,0.6t$) for cycle 1,2, and 3 respectively. One should expect that pumping can only happen for cycle $2$ (green continuous line) as it winds around the topological phase transition point $V=0$. To study the TCP, we calculate a quantity called polarization given by
@@
where $i_{0}=(L+1)/2$ for a system of $L=200$ sites for all the three cycles. Here $\psi(\tau)$ represents the ground state wave function of the Hamiltonian defined in Eq. (18). The amount of charge pumped ($Q$) in a cycle is given as
$Q=\int_{0}^{1}d\tau~\partial_{\tau}P(\tau).$ (20)
Fig. 13 clearly demonstrates a smooth change in the polarization from $0.5$ to 0.5-0.5, indicating a total charge transfer of $|Q|=1$ for cycle 2. This signifies robust and quantized pumping occurring in cycle 2 as it encloses the gap-closing topological phase transition point as mentioned earlier. In contrast, cycle 1 exhibits a breakdown of pumping (due to the appearance of a discontinuity in the polarization), while cycle 3 shows no pumping at all Hayward et al. (2018). This analysis complements the topological phase transition between the two BO phases shown in the phase diagram [Fig. 2(a)].